# Auditing Black-Box LLM APIs with a Rank-Based Uniformity Test

### v3

@@ 5.1 Experimental Setup
Gemma–4bit 0.392 0.214 0.017
Gemma–8bit 0.049 0.043 0.001
Llama–4bit 0.642 0.625 0.474
Llama–8bit 0.132 0.158 0.005
Mistral–4bit 0.586 0.500 0.330
Mistral–8bit 0.049 0.090 0.006
@@
Mistral Dan 0.895 0.802 0.873
Anti-Dan 0.893 0.781 0.872
Evil-Bot 0.892 0.766 0.873
Gemma Dan 0.888 0.757 0.867
Anti-Dan 0.858 0.816 0.854
Evil-Bot 0.893 0.753 0.871
Table 2: Side-by-side comparison of statistical-power AUCs. Left: quantized variants; right: jail-breaking prompts.
5.2 Detecting Quantization
We consider the setting where the API provider uses a quantized variant to substitute the claimed model. We evaluate three detection methods on quantized variants of Llama-3.2-3B-Instruct [^1] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct, Mistral-7B-Instruct-v0.3222https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, and Gemma-2-9B-it [^3] https://huggingface.co/google/gemma-2-9b-it, comparing each model to its 4-bit and 8-bit quantized counterparts. As shown in Table 2(a), none of the methods succeed in reliably detecting substitution for the 8-bit variants of Gemma and Mistral, where statistical power AUC remains near zero across the board. In the remaining four settings, RUT outperforms MMD and the KS baseline in 3 out of the 4 cases, demonstrating superior sensitivity to quantization-induced distributional shifts. Full statistical power curves for AUCs are provided in Appendix B.1.
@@ 5.3 Detecting Jailbreaks
We consider the setting where the API provider secretly appends a hidden jailbreaking system prompt to user queries. To evaluate this scenario, we use two base models: Mistral-7B-Instruct-v0.3 and Gemma-2-9B-it. For each model, we construct a test using three representative jailbreaking prompts Dan, Anti-Dan, and Evil-Bot adapted from Shen et al. (2024).
As shown in Table 2(b), all six jailbreak cases are reliably detected, with statistical power AUC values consistently above 0.75. RUT achieves the highest power in all 6 settings, demonstrating its superior sensitivity to model deviations caused by hidden jailbreaking prompts. Full statistical power curves for AUCs are provided in Appendix B.2.
@@ 5.4 Detecting SFT
Figure 3: Statistical power AUC across epochs for detecting SFT model substitution.
@@
As shown in Figure 3, RUT consistently achieves higher statistical power AUC than both the KS and MMD baselines across all fine-tuning configurations. Notably, our method detects behavioral changes within the first epoch of fine-tuning, demonstrating strong sensitivity to early-stage distributional shifts. While all methods improve with additional training, RUT remains the most robust across both models and datasets. Full statistical power curves for AUCs are provided in Appendix B.3.
@@ 5.5 Detecting Full Model Replacement
Figure 4: Statistical power AUC for detecting full model replacement. Each cell shows the AUC score between a reference and a target model. Diagonal values represent self-comparisons.
@@ 5.6 Detecting Real API Providers
Table 3: Statistical power for detecting differences from the baseline model deployed on an A6000 GPU. A100 denotes the same model run locally on an A100 GPU; other entries are actual API providers. Values $>0.5$ indicate significant behavioral deviation. Green = no significant difference; Red = significant difference.
@@
We evaluate our detection methods on three base models—Llama-3.2-3B-Instruct, Mistral-7B-Instruct-v0.3, and Gemma-2-9B-it—each deployed through multiple API providers. Local inference on an A100 GPU serves as the baseline. As shown in Table 3, all tests correctly identify behavioral equivalence in local deployments.
Across all settings, RUT and MMD generally agree in detecting significant deviations across providers, offering mutual validation for their behavioral sensitivity. The KS test exhibits similar trends but with notably lower sensitivity.
An exception arises in the Mistral + HF Inference setting, where MMD yields a power of 1.000 while other tests are below 0.2. Upon investigation, we suspect that the discrepancy is due to a tokenization mismatch: the Hugging Face Inference API consistently omits the leading whitespace present in the reference outputs. Because MMD uses character-level Hamming distance, this formatting difference inflates the score. After restoring the missing space, the MMD score drops to 0.211, aligning with other tests. This illustrates RUTt’s robustness to minor decoding mismatches that can mislead string-based metrics.
@@ 6 Conclusion
Using prompts from the WildChat dataset (Zhao et al., 2024) for testing, we aim to avoid detection of our testing attempts. However, we have not empirically validated its effectiveness at evading detection, especially against detection methods based on the correlation across prompts. In future work, examining the method’s detectability and improving it with adaptive methods for prompt selection shall be a priority. Also, our test requires a locally deployed authentic model, limiting its capability to test black-box models.
@@ A.1 AUROC Algorithm
Input: Prompt set $\mathcal{D}$; models ${\pi}_{\text{ref}}$, ${\pi}_{\text{tgt}}$; decoding parameters $\varphi=(\tau,L)$, where $\tau$ is temperature and $L$ is the maximum generation length; number of prompts $n$; completions per prompt per model $m$; score functions $\{\delta_{1},\dots,\delta_{K}\}$.
@@ 8       $L_{i}\leftarrow\{0\}^{m}\cup\{1\}^{m}$;
10      for $\delta\in\{\delta_{1},\dots,\delta_{K}\}$ do
@@ 12             Store $A_{i}^{\delta}\leftarrow\mathrm{AUROC}(S_{i},L_{i})$;
16for $\delta\in\{\delta_{1},\dots,\delta_{K}\}$ do

### v4

@@ 5.2 Detecting Quantization
Gemma–4bit 0.392 $84^{-}$$103^{+}$ 0.214 $10^{-}11$$12^{+}$ 0.017 $32^{-}$$34^{+}$
Gemma–8bit 0.049 $^{-}59+59^{+}$ 0.043 $62^{-}$$64^{+}$ 0.001 $08^{-}$$10^{+}$
Llama–4bit 0.642 $84^{-}$$82^{+}$ 0.625 $89^{-}$$93^{+}$ 0.474 $88^{-}$$89^{+}$
Llama–8bit 0.132 $82^{-}$$81^{+}$ 0.158 $110^{-}$$109^{+}$ 0.005 $17^{-}$$19^{+}$
Mistral–4bit 0.586 $95^{-}$$94^{+}$ 0.500 $92^{-}$$97^{+}$ 0.330 $100^{-}$$102^{+}$
Mistral–8bit 0.049 $56^{-}$$61^{+}$ 0.090 $77^{-}$$84^{+}$ 0.006 $0^{-}22$$2^{+}$
@@
Mistral Dan 0.895 $46^{-}$$45^{+}$ 0.802 $63^{-}$$55^{+}$ 0.873 $37^{-}$$36^{+}$
Anti-Dan 0.893 $43^{-}$$41^{+}$ 0.781 $2^{-}55$$5^{+}$ 0.872 $46^{-}$$49^{+}$
Evil-Bot 0.892 $^{-}44+44^{+}$ 0.766 $56^{-}$$59^{+}$ 0.873 $^{-}37+37^{+}$
Gemma Dan 0.888 $3^{-}44$$4^{+}$ 0.757 $65^{-}$$68^{+}$ 0.867 $34^{-}$$36^{+}$
Anti-Dan 0.858 $49^{-}$$53^{+}$ 0.816 $60^{-}$$52^{+}$ 0.854 $^{-}31+31^{+}$
Evil-Bot 0.893 $43^{-}$$42^{+}$ 0.753 $65^{-}$$63^{+}$ 0.871 $^{-}36+36^{+}$
Table 2: Statistical power AUCs. All confidence interval ranges are reported in $\times 10^{-4}$
@@ 5.3 Detecting Jailbreaks
We consider the setting where the API provider secretly appends a hidden jailbreaking system prompt to user queries. To evaluate this scenario, we use two base models: Mistral-7B-Instruct-v0.3 and Gemma-2-9B-it. For each model, we construct a test using three representative jailbreaking prompts Dan, Anti-Dan, and Evil-Bot adapted from Shen et al. (2024). As shown in Table 2(b), all jailbreak cases are reliably detected, with power AUCs consistently above 0.75. RUT achieves the highest power in all 6 settings, demonstrating its superior sensitivity to model deviations caused by hidden jailbreaking prompts. Full statistical power curves for AUCs are provided in Appendix B.2.
@@ 5.4 Detecting SFT
As shown in Figure 3, RUT consistently achieves higher power AUC than both the KS and MMD baselines across all fine-tuning configurations. Notably, our method detects behavioral changes within the first epoch of fine-tuning, demonstrating strong sensitivity to early-stage distributional shifts. While all methods improve with additional training, RUT remains the most robust across both models and datasets. Full statistical power curves for AUCs are provided in Appendix B.3.
Figure 3: AUC of SFT checkpoints across epochs.
@@ 5.5 Detecting Full Model Replacement
Figure 4: AUC for detecting full model replacement. Each cell shows the AUC score between a reference and a target model. Diagonal values represent self-comparisons.
@@
5.6 Detecting Query Domains
We now evaluate the robustness of RUT across more query domains. Beyond WildChat which reflects general conversational traffic, we consider two specialized datasets: BigCodeBench (Zhuo et al., 2024) for programming tasks and MATH (Hendrycks et al., 2021) for mathematical problem solving. We adopt the quantization setup in Section 5.2 as this setting is both challenging and practically relevant. See full power curves in Appendix B.5.
As shown in Table 3, the overall detection powers mirror those observed in Section 5.2. Detection remains difficult for 8-bit quantized Gemma and Mistral, where all methods fail to achieve meaningful power. RUT consistently shows high detectability in the remaining cases, outperforming MMD and KS in seven out of the eight cases. These results show that RUT remains effective in both math and code domains, reinforcing its generalizability to diverse query distributions.
Table 3: Statistical power AUC for detecting quantized variants across query domains. Bold = best method; gray = none reliable. All confidence interval ranges are reported in $\times 10^{-4}$
(a) BigCodeBench (Zhuo et al., 2024)
Model RUT MMD KS
Gemma–4bit 0.136 $98^{-}$$95^{+}$ 0.170 $89^{-}$$103^{+}$ 0.000 $22^{-}$$12^{+}$
Gemma–8bit 0.048 $62^{-}$$71^{+}$ 0.052 $69^{-}$$58^{+}$ 0.000 $05^{-}$$09^{+}$
Llama–4bit 0.593 $54^{-}$$50^{+}$ 0.353 $81^{-}$$80^{+}$ 0.524 $48^{-}$$50^{+}$
Llama–8bit 0.093 $4^{-}33$$3^{+}$ 0.143 $56^{-}$$58^{+}$ 0.021 $12^{-}$$23^{+}$
Mistral–4bit 0.462 $76^{-}$$84^{+}$ 0.326 $91^{-}$$88^{+}$ 0.183 $43^{-}$$51^{+}$
Mistral–8bit 0.041 $73^{-}$$74^{+}$ 0.058 $67^{-}$$69^{+}$ 0.011 $77^{-}$$83^{+}$
(b) Math (Hendrycks et al., 2021)
Model RUT MMD KS
Gemma–4bit 0.297 $102^{-}$$133^{+}$ 0.237 $99^{-}$$105^{+}$ 0.059 $76^{-}$$75^{+}$
Gemma–8bit 0.048 $^{-}64+64^{+}$ 0.067 $72^{-}$$75^{+}$ 0.002 $33^{-}$$35^{+}$
Llama–4bit 0.532 $99^{-}$$94^{+}$ 0.468 $101^{-}$$85^{+}$ 0.412 $5^{-}44$$4^{+}$
Llama–8bit 0.253 $72^{-}$$73^{+}$ 0.177 $5^{-}66$$6^{+}$ 0.118 $75^{-}$$80^{+}$
Mistral–4bit 0.360 $33^{-}$$37^{+}$ 0.216 $66^{-}$$65^{+}$ 0.160 $33^{-}$$34^{+}$
Mistral–8bit 0.059 $44^{-}$$47^{+}$ 0.060 $^{-}58+58^{+}$ 0.019 $45^{-}$$43^{+}$
@@ 5.7 Detecting Real API Providers
Table 4: Statistical power for detecting differences from the model deployed on an A6000 GPU. A100 denotes the same model on an A100 GPU. Values >0.5>0.5 indicate significant behavioral deviation. Green = no significant difference; Red = significant difference.
@@
We evaluate three models—Llama-3.2-3B, Mistral-7B, and Gemma-2-9B—across multiple API providers, using local A100 inference as the baseline. As shown in Table 4, all tests correctly confirm behavioral equivalence in local deployments.
Across all settings, RUT and MMD generally agree in detecting significant deviations across providers, offering mutual validation for their behavioral sensitivity. The KS test exhibits similar trends but with notably lower sensitivity. One exception is Mistral + HF Inference, where MMD yields a power of 1.0 while others remain below 0.2. We trace this to a tokenization mismatch: the HF Inference API consistently omits the leading whitespace present in the reference outputs. Since MMD uses string Hamming distance, the formatting difference inflates the score. After restoring the missing space, the MMD score drops to 0.211, aligning with other tests. This shows the robustness of RUT to minor decoding mismatches that can mislead string-based metrics.
@@ 6 Conclusion
We have not empirically validated effectiveness of RUT against an adversary with full knowledge of the auditing method. For example, an attacker could selectively reroute prompts that are expected to produce atypical log-rank statistics. Assessing the method’s robustness against more powerful adversaries is an important next step. In addition, RUT requires access to a locally deployed authentic reference model, which limits its applicability to open-sourced models. Exploring ways to relax this requirement would broaden the method’s applicability.
@@
Acknowledgements
Thanks Muru Zhang and Oliver Liu for valuable discussions. The work was supported in part by the National Science Foundation (NSF) under Award No. 2427856. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of NSF.
@@ A.1 AUROC Algorithm
Input: Prompt set $\mathcal{D}$; models ${\pi}_{\text{ref}}$, ${\pi}_{\text{tgt}}$; decoding parameters $\varphi=(\tau,L)$, where $\tau$ is temperature and $L$ is the maximum generation length; number of prompts $n$; number of completions per prompt per model $m$; score functions $\{\varphi_{1},\dots,\varphi_{K}\}$.
@@ 8 $L_{i}\leftarrow\{0\}^{m}\cup\{1\}^{m}$;
10 for $\delta\in\{\varphi_{1},\dots,\varphi_{K}\}$ do
@@ 12      Store $A_{i}^{\delta}\leftarrow\mathrm{AUROC}(S_{i},L_{i})$;
16for $\delta\in\{\varphi_{1},\dots,\varphi_{K}\}$ do
@@ B.4 Full Statistic Power Curves for Detecting Full Model Replacement
B.5 Full Statistic Power Curves for Detecting More Query Domains
We present the full statistical power curves, showing the relationship between substitution rate and detection power, corresponding to the experiments on detecting quantized model in extra domains described in Section 5.6. These curves are used to compute the power AUC values reported in the main paper and illustrate each method’s detection power across different levels of substitution.
B.6 Case Study on Detecting Decoding Parameters
(a) Gemma-2-9B-it on MATH
(b) Llama-3.2-3B-Instruct on MATH
(c) Gemma-2-9B-it on WildChat
(d) Llama-3.2-3B-Instruct on WildChat
Figure 5: Statistical power AUC for detecting decoding parameter mismatches (temperature, top-$p$) across models and datasets. Each cell compares outputs under a specific decoding configuration against the default $(0.5,1.0)$; higher values indicate stronger detectability.
Setup.
As a case study, we test whether detection methods can identify changes in decoding parameters, focusing on sampling temperature and top-$p$ for nucleus sampling (Holtzman et al., 2020). We compare responses generated under different parameter settings against the default configuration of temperature=0.5 and top-pp=1.
Findings.
We perform the experiments across Gemma-2-9B-it and Llama-3.2-3B-Instruct on MATH (Hendrycks et al., 2021) and WildChat (Zhao et al., 2024). Based on the results in Figure 5, RUT achieves consistently higher detection power across decoding configurations compared to MMD and KS. This sensitivity is desirable in practice, since when the API providers expose decoding controls to users, a reliable detection method should be able to flag deviations arising not only from model substitution but also from anomalous decoding configurations.
Appendix C Theroretical Analysis of RUT
In this section, we provide a formal analysis to characterize the statistical power of RUT and its sample efficiency under principled assumptions about language model output distributions and adversarial perturbations.
We model two basic attack types: (1) reparameterization attack, which preserve ranking of tokens but smoothly alter token probabilities via a small parameter shift $s\mapsto s+\epsilon$; and (2) reordering attack, which randomly samples a token $t$ according to the distribution, and then swaps $t$ with the highest-probability token, thereby making $t$ the new top-1 token. Real-world attacks almost always combine both, and tend to easier to detect than either alone.
Throughout the proof, we assume that RUT uses the log-rank score function. Based on the famous Zip’s law in statistical linguistics (Piantadosi, 2014), we assume that the output token distributions at any position follows a Zeta distribution with parameter $s$, where $s\in(1,+\infty)$ reflects the heavy-tailedness of the distribution.
It can be easily shown that the FPR of RUT is $\alpha$, the significance level that we choose. Below, we focus on type II error: the probability of missing an attack when one exists.
Lemma C.1.
Let the rank of a token be a random variable $K$ on $\mathbb{Z}^{+}$ with probability mass function $p(k)=k^{-\alpha}/\zeta(\alpha)$ for $\alpha>1$. Let $x=\log k$. The survival function $S(x)=P(\log K\geq x)$, for large $k=e^{x}$, has the asymptotic form
$S(x)=\frac{e^{(1-\alpha)x}}{\zeta(\alpha)(\alpha-1)}(1+o(1))$
Proof.
The survival function is the probability that the rank $K$ is greater than or equal to some value $m$. By definition, this is the sum of the probabilities of all ranks from $m$ to infinity.
$S(\log m)=P(K\geq m)=\sum_{k=m}^{\infty}p(k)=\frac{1}{\zeta(\alpha)}\sum_{k=m}^{\infty}k^{-\alpha}$
For large $m$, this sum can be approximated by its corresponding integral. The leading term of the Euler-Maclaurin formula shows that the sum and integral are asymptotically equivalent.
$\sum_{k=m}^{\infty}k^{-\alpha}$ $=\int_{m}^{\infty}t^{-\alpha}dt+O(m^{-\alpha})$
$=\left[\frac{t^{-\alpha+1}}{1-\alpha}\right]_{m}^{\infty}+O(m^{-\alpha})$
$=0-\frac{m^{1-\alpha}}{1-\alpha}+O(m^{-\alpha})$
$=\frac{m^{1-\alpha}}{\alpha-1}+O(m^{-\alpha})$
$=\frac{m^{1-\alpha}}{\alpha-1}(1+O(m^{-1}))$
$=\frac{m^{1-\alpha}}{\alpha-1}(1+o(1))$
Substituting this result back into the expression for the survival function, we obtain
$S(\log m)=\frac{m^{1-\alpha}}{\zeta(\alpha)(\alpha-1)}(1+o(1))$
The analysis is performed on the distribution of the log-rank score, $x=\log k$. We therefore let $x=\log m$, which implies $m=e^{x}$. Substituting this into the expression above yields the final asymptotic form for the survival function of the log-rank.
$S(x)=\frac{(e^{x})^{1-\alpha}}{\zeta(\alpha)(\alpha-1)}(1+o(1))=\frac{e^{(1-\alpha)x}}{\zeta(\alpha)(\alpha-1)}(1+o(1))$
For the remainder of this analysis, we denote this survival function, which serves as the CDF for our rank-based test, by $F(x,\alpha)$. ∎
Lemma C.2.
For a hypothesis test using the Cramér-von Mises (CvM) statistic to distinguish a null hypothesis $H_{0}$ from an alternative hypothesis $H_{1}$ with a fixed significance level $\alpha_{err}$ and a fixed statistical power $1-\beta_{err}$, the required sample size $n$ is assymptotically inversely proportional to the CvM distance $\omega_{\infty}^{2}$ between the two distributions, i.e.
$n=\Theta\left(\frac{1}{\omega_{\infty}^{2}}\right),\text{ as }\omega^{2}_{+\infty}\to 0\text{,}$
where the constant coefficient is given by (1).
Proof.
Let the ranks $r_{i}$ for $i=1,\dots,n$ be drawn from a distribution with CDF $F(x)$. The hypothesis test is $H_{0}:F(x)=U(x)$ versus $H_{1}:F(x)=F_{1}(x)$, where $U(x)=x$ is the CDF of the Uniform[0,1] distribution. The CvM test statistic is
$\omega^{2}=n\int_{0}^{1}(F_{n}(x)-U(x))^{2}dx$
where $F_{n}(x)$ is the empirical CDF. The test rejects $H_{0}$ if the observed statistic $\omega^{2}$ exceeds a critical value $c_{\alpha}$. The power of the test is $1-\beta_{err}=P_{H_{1}}(\omega^{2}>c_{\alpha})$. The analysis of this power requires the distribution of $\omega^{2}$ under $H_{1}$.
While the distribution of $\omega^{2}$ under $H_{0}$ is a non-normal, weighted sum of chi-squared variables, its distribution under a fixed alternative $H_{1}$ is asymptotically normal as $n\to\infty$. As a standard result in statistics, this follows from the central limit theorems of U-statistics, which establish that the standardized statistic converges in distribution to a standard normal variable:
$\frac{\omega^{2}-E_{H_{1}}[\omega^{2}]}{\sqrt{\text{Var}_{H_{1}}(\omega^{2})}}\xrightarrow{d}N(0,1)$
The expected value $E_{H_{1}}[\omega^{2}]$ contains a non-centrality parameter that grows linearly with $n$, such that $E_{H_{1}}[\omega^{2}]=n\cdot\omega_{\infty}^{2}+O(1)$, where the CvM distance $\omega_{\infty}^{2}$ is defined as
$\omega_{\infty}^{2}=\int_{0}^{1}(F_{1}(x)-U(x))^{2}dx$
The standard deviation, $\sqrt{Var_{H_{1}}(\omega^{2})}$, grows slower than the mean, and the standardized deviation $\sigma_{1}=\sqrt{\text{Var}_{H_{1}}(\omega^{2})}/n$ is asymptotically $O(1)$. We then analyze the power for large $n$:
$1-\beta_{err}=P\left(Z>\frac{c_{\alpha}-(n\cdot\omega_{\infty}^{2}+O(1))}{\sigma_{1}}\right)$
where $Z$ is a standard normal variable. For the power to be a desired constant, the argument to the probability function must also be a constant, denoted $z_{\beta}$.
$\frac{c_{\alpha}-n\cdot\omega_{\infty}^{2}+O(1)}{\sigma_{1}}=-z_{\beta}$
Solving this equation for $n$ gives
$n\cdot\omega_{\infty}^{2}=c_{\alpha}+z_{\beta}\sigma_{1}+O(1)=\Theta(1)\quad\quad(1)$
Since the right-hand side consists of terms that are constant for a fixed significance level and power, it follows that $n\cdot\omega_{\infty}^{2}$ must be constant. This implies the inverse proportionality $n\propto 1/\omega_{\infty}^{2}$. ∎
Theorem C.1.
Consider a reparameterization attack that perturbs the token rank distribution’s Zipfian parameter $\alpha$ by an amount $\epsilon$. For RUT to achieve a constant Type II error rate:
  • When $\alpha\to 1^{+}$ (moderate heavy-tailedness), a target sample size of $n=\Theta(\epsilon^{-2}(\alpha-1)^{2})$ is sufficient.
  • When $\alpha\to+\infty$ (extreme heavy-tailedness), a target sample size of $n=\Theta(\epsilon^{-2}\alpha^{5})$ is sufficient.
Proof.
A reparameterization attack shifts the distribution parameter from $\alpha$ to $\alpha^{\prime}=\alpha+\epsilon$. The CvM distance $\omega_{\infty}^{2}$ is proportional to $\epsilon^{2}$ and the integral of the squared derivative of the log-rank CDF, $F(x,\alpha)$, with respect to $\alpha$. First, we derive the asymptotic form of this derivative. Let $F(x,\alpha)=e^{(1-\alpha)x}D(\alpha)^{-1}$, where $D(\alpha)=\zeta(\alpha)(\alpha-1)$. The partial derivative is
$\frac{\partial F(x,\alpha)}{\partial\alpha}$ $=\frac{\partial}{\partial\alpha}\left(e^{(1-\alpha)x}D(\alpha)^{-1}\right)$
$=-xe^{(1-\alpha)x}D(\alpha)^{-1}-e^{(1-\alpha)x}D(\alpha)^{-2}D^{\prime}(\alpha)$
$=-e^{(1-\alpha)x}\left(\frac{x}{D(\alpha)}+\frac{D^{\prime}(\alpha)}{D(\alpha)^{2}}\right)$
where
$D^{\prime}(\alpha)=\frac{d}{d\alpha}(\zeta(\alpha)(\alpha-1))=\zeta^{\prime}(\alpha)(\alpha-1)+\zeta(\alpha)$
As $\alpha\to\infty$, we have the known asymptotic behaviors $\zeta(\alpha)=1+o(1)$ and $\zeta^{\prime}(\alpha)=o(1)$. Therefore,
$D(\alpha)=\zeta(\alpha)(\alpha-1)=(1+o(1))(\alpha-1)=(\alpha-1)(1+o(1))$
$D^{\prime}(\alpha)=\zeta^{\prime}(\alpha)(\alpha-1)+\zeta(\alpha)=o(1)(\alpha-1)+(1+o(1))=1+o(1)$
Substituting these into the expression for the derivative gives
$\frac{\partial F(x,\alpha)}{\partial\alpha}$ $=-e^{(1-\alpha)x}\left(\frac{x}{(\alpha-1)(1+o(1))}+\frac{1+o(1)}{(\alpha-1)^{2}(1+o(1))^{2}}\right)$
$=-e^{(1-\alpha)x}\left(\frac{x}{\alpha-1}(1+o(1))+\frac{1}{(\alpha-1)^{2}}(1+o(1))\right)$
$=-e^{(1-\alpha)x}\Theta\left(\alpha^{-1}x+\alpha^{-2}\right)$
The CvM distance $\omega_{\infty}^{2}$ is proportional to
$\epsilon^{2}\int_{0}^{\infty}(\frac{\partial F}{\partial\alpha})^{2}p(x)dx$
The probability density function is $p(x)=-\frac{\partial F(x,\alpha)}{\partial x}=\frac{e^{(1-\alpha)x}}{\zeta(\alpha)}$.
$\omega_{\infty}^{2}$ $=\Theta\left(\epsilon^{2}\int_{0}^{\infty}\left(-e^{(1-\alpha)x}(\alpha^{-1}x+\alpha^{-2})\right)^{2}\frac{e^{(1-\alpha)x}}{\zeta(\alpha)}dx\right)\quad\quad(2)$
$=\Theta\left(\frac{\epsilon^{2}}{\zeta(\alpha)}\int_{0}^{\infty}e^{2(1-\alpha)x}\left(\alpha^{-2}x^{2}+2\alpha^{-3}x+\alpha^{-4}\right)e^{(1-\alpha)x}dx\right)$
$=\Theta\left(\frac{\epsilon^{2}}{\zeta(\alpha)}\int_{0}^{\infty}e^{3(1-\alpha)x}(\alpha^{-2}x^{2}+O(\alpha^{-3}x)+\alpha^{-4})dx\right)$
We now analyze this integral expression in two asymptotic regimes. As $\alpha\to 1^{+}$, let $k=3(\alpha-1)\to 0^{+}$. The integral becomes $\int_{0}^{\infty}e^{-kx}(\alpha^{-2}x^{2}+O(\alpha^{-3}x)+\alpha^{-4})dx$. We use the standard identity $\int_{0}^{\infty}e^{-kt}t^{n}dt=n!/k^{n+1}$.
Integral $=\alpha^{-2}\int_{0}^{\infty}e^{-kx}x^{2}dx+\alpha^{-4}\int_{0}^{\infty}e^{-kx}dx+O\left(\int_{0}^{\infty}e^{-kx}xdx\right)$
$=\alpha^{-2}\frac{2!}{k^{3}}+\alpha^{-4}\frac{0!}{k^{1}}+O(k^{-2})$
$=\frac{2\alpha^{-2}}{(3(\alpha-1))^{3}}+\frac{\alpha^{-4}}{3(\alpha-1)}+O((\alpha-1)^{-2})\quad\quad(3)$
As $\alpha\to 1^{+}$, the dominant term is $\Theta((\alpha-1)^{-3})$. Since $\zeta(\alpha)^{-1}=\Theta(\alpha-1)$, the CvM distance is
$\omega_{\infty}^{2}=\Theta\left(\epsilon^{2}\cdot(\alpha-1)\cdot(\alpha-1)^{-3}\right)=\Theta\left(\epsilon^{2}(\alpha-1)^{-2}\right)$
By Lemma C.2, the required sample size is $n=\Theta(\epsilon^{-2}(\alpha-1)^{2})$.
As $\alpha\to+\infty$, a re-calculation results in $(2)$ being of order $\Theta(\alpha^{-5})$; note that $(3)$ no longer apply due to change in the limiting condition. Since $\zeta(\alpha)=1+o(1)$, we have
$\omega_{\infty}^{2}=\Theta\left(\epsilon^{2}\cdot\alpha^{-5}\right)$
By Lemma C.2, the sample size is $n=\Theta\left(\epsilon^{-2}\alpha^{5}\right)$. ∎
Theorem C.2.
Consider a reordering attack where a sampled token is made the new top-1 token. For RUT to achieve a constant Type II error rate:
  • When $\alpha\to 1^{+}$ (moderate heavy-tailedness), a target sample size of $n=\Theta(1)$ is sufficient.
  • When $\alpha\to+\infty$ (extreme heavy-tailedness), a target sample size of $n=\Theta(2^{\alpha})$ is sufficient.
Proof.
For a reordering attack, the CvM distance $\omega_{\infty}^{2}$ is the expected squared difference in the log-rank scores, given by
$\omega^{2}_{\infty}=\sum_{k=1}^{+\infty}\frac{k^{-\alpha}(1-k^{-\alpha})\log k}{\zeta(\alpha)^{2}}$
Expanding the numerator and splitting the expression into two sums yields
$\omega^{2}_{\infty}=\frac{1}{\zeta(\alpha)^{2}}\left(\sum_{k=1}^{\infty}\frac{\log k}{k^{\alpha}}-\sum_{k=1}^{\infty}\frac{\log k}{k^{2\alpha}}\right)$
Using the identity $\zeta^{\prime}(s)=-\sum_{k=1}^{\infty}k^{-s}\log k$, we obtain
$\omega^{2}_{\infty}=\frac{1}{\zeta(\alpha)^{2}}\left((-\zeta^{\prime}(\alpha))-(-\zeta^{\prime}(2\alpha))\right)=\frac{\zeta^{\prime}(2\alpha)-\zeta^{\prime}(\alpha)}{\zeta(\alpha)^{2}}$
We analyze this closed-form expression. As $\alpha\to 1^{+}$, we use the Laurent series expansions $\zeta(\alpha)=(\alpha-1)^{-1}+\gamma+O(\alpha-1)$ and $\zeta^{\prime}(\alpha)=-(\alpha-1)^{-2}+\gamma_{1}+O(\alpha-1)$, where $\gamma$ and $\gamma_{1}$ are Stieltjes constants. The term $\zeta^{\prime}(2\alpha)$ converges to the finite constant $\zeta^{\prime}(2)$. The numerator is dominated by $\zeta^{\prime}(\alpha)$, so
$\omega_{\infty}^{2}=\frac{-(\alpha-1)^{-2}+O(1)}{((\alpha-1)^{-1}+O(1))^{2}}=\frac{-(\alpha-1)^{-2}(1+O((\alpha-1)^{2}))}{(\alpha-1)^{-2}(1+O(\alpha-1))^{2}}=1+O(\alpha-1)=\Theta(1)$
By Lemma C.2, a sample size of $n=\Theta(1)$ is sufficient.
As $\alpha\to+\infty$, we have $\zeta(\alpha)=1+o(1)$. The derivative is dominated by its leading term from the series expansion, $\zeta^{\prime}(\alpha)=-2^{-\alpha}\ln 2(1+o(1))$. The term $\zeta^{\prime}(2\alpha)=-4^{-\alpha}\ln 2(1+o(1))$ is of a lower order. Therefore,
$\omega_{\infty}^{2}$ $=\frac{-\zeta^{\prime}(\alpha)(1-\zeta^{\prime}(2\alpha)/\zeta^{\prime}(\alpha))}{(\zeta(\alpha))^{2}}$
$=\frac{-(-2^{-\alpha}\ln 2(1+o(1)))(1-\frac{-4^{-\alpha}\ln 2(1+o(1))}{-2^{-\alpha}\ln 2(1+o(1))})}{(1+o(1))^{2}}$
$=\frac{2^{-\alpha}\ln 2(1+o(1))(1-2^{-\alpha}(1+o(1)))}{(1+o(1))^{2}}$
$=2^{-\alpha}\ln 2(1+o(1))=\Theta(2^{-\alpha})$
By Lemma C.2, the required sample size is $n=\Theta(2^{\alpha})$. ∎